# The Dark Side of Function Calling: Pathways to Jailbreaking Large Language Models

### v3

@@ 2 Related Work
Jailbreak Risks in Function Calling. To the best of our knowledge, no current research has revealed the jailbreak risks associated with function calling. The closest attempt at a successful jailbreak is by Pelrine et al. (2023). They designed three different functions to try to bypass the safety alignment of the GPT-4 model, but all of these attempts failed. Upon analysis, we found that the reason for this failure lies in their approach; they used harmful behavior as the argument for the functions and attempted to extract the jailbreak content in the final response. However, as our work shown, vulnerabilities often lie hidden in the generation of the argument itself.
@@ 4.2 Analyzing Why Function Calls Cause Jailbreaks (RQ 2)
Table 3: ASR for different defensive prompt insertion positions.

### v4

@@ 2 Related Work
Jailbreak Risks in Function Calling. To the best of our knowledge, no current research has revealed the jailbreak risks associated with function calling. The closest attempt at a successful jailbreak is by Pelrine et al. (2023). They designed three different functions to try to bypass the safety alignment of the GPT-4 model, but all of these attempts failed. The reason for this failure lies in their approach; they used harmful behavior as the argument for the functions and attempted to extract the jailbreak content in the final response. However, as our work shown, vulnerabilities often lie hidden in the generation of the argument itself.
@@ 4.2 Analyzing Why Function Calls Cause Jailbreaks (RQ 2)
Table 3: ASR for different defensive prompt insertion positions.
@@ 5 Conclusion
Limitations
The primary limitation of this study is its narrow testing scope. We focused on designing and evaluating a single specific jailbreak function, which does not account for the wide range of potential attack variations. As technology evolves, attackers are likely to create more sophisticated and innovative jailbreak functions (in fact, variants of our work have already emerged in the community [^1] https://www.reddit.com/r/ChatGPTJailbreak/comments/1eq65ng/function_call_jailbreak/), which may outpace our current defense strategies. Therefore, there is a pressing need to develop more robust and adaptable defense mechanisms to effectively counter future jailbreak function attacks.
Ethical Considerations
Before LLMs become more intelligent and deeply integrated into everyday life, our work has timely revealed the jailbreak risks associated with function calling. In the short term, this might raise the immediate risk of jailbreak attacks, but in the long run, our findings serve as a crucial reminder for LLM researchers and developers to prioritize the security of all model interaction methods, including function calling. By highlighting these risks early on, we aim to encourage the development of more comprehensive and robust safety measures that will ensure the secure and responsible use of LLMs.
Acknowledgements
This work was supported in part by the National Key R&D Program of China (2023YFB3107505), in part by Shaanxi Natural Science Funds for Distinguished Young Scholars (2023-JC-JQ-52), in part by the Natural Science Foundation of China (62302371), in part by the Postdoctoral Fellowship Program of CPSF (GZC20232035), and in part by the Fundamental Research Funds for the Central Universities (ZYTS24083).